OpenStudy (anonymous):

The product of the roots of 8x² - 2x = 1 is

ganeshie8 (ganeshie8):

write it in standard form

OpenStudy (nurali):

8x^2-2x-1=0 8x^2-4x+2x-1=0 4x(2x-1)+1(2x-1)=0 (4x+1)(2x-1)=0 x=-1/4 & 1/2 -1/4 * 1/2=-1/8